A Bibliography of The Journal of Visualization and Computer Animation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.17

Title word cross-reference

3 [222, 131, 44, 120, 217, 151, 168, 219, 78, 214, 188, 189, 114, 47, 164, 40, 10, 190, 191, 158, 16]. N [89].

-body [89]. -D [16].

3-D [272]. 3D [238, 239, 246, 247, 281, 275, 251, 276].

’93 [75]. ’95 [105, 121]. ’97 [142].

ability [161]. Abstraction [84]. Acceleration [135, 194]. Access [110]. Actor [79]. actors [23]. Adaptation [140]. Adapting [289]. Adaptive [265, 272]. Advanced [227, 57]. Ageing [92]. agents [78, 262]. aided [156, 137]. Algorithm [5, 152, 137, 171, 44, 280, 48, 215, 218, 54, 251, 283]. algorithms [230]. ambition [278]. Analysis [104, 228, 109, 54, 13]. anatomical [238, 239]. ancient [281]. animal [200]. Animated [132, 78, 55, 118, 214, 225, 260]. Animating [212, 116, 30, 1, 27]. Animation [61, 82, 120, 64, 8, 157, 76, 7, 60, 186, 152, 39, 2, 92, 67, 221, 201, 210, 21, 259, 24, 261, 215, 173, 203, 17, 159, 199, 250, 19, 178, 34, 216, 184, 13, 4, 45, 26, 177, 15, 231, 16]. Animations [104]. anisotropic [249]. annotated [11, 14]. Antarctic [112]. anthropomorphic [190]. Application [290, 48, 219, 181, 158]. applications [277, 168, 275]. Applied [69]. Applying [83]. Approach [151, 87, 122, 86, 39, 206, 174, 167, 199, 49, 284, 4, 16].

Approximation [67, 228]. Arbitrary [118]. Architectural [150]. architecture [23]. archival [274]. Arm [137, 288]. Arrangement [102]. art [277, 244, 245]. art-oriented [277]. Articulated [88, 71, 27]. Arts [37]. ArtShop [277]. assisted [129]. attention [261]. Augmenting [256]. aurora [268]. autistic [286]. automata [168]. Automated [29, 274]. Automatic [129, 79]. automaton [158]. automaton-based [158]. Autumn [102]. Avoidance [56].

B [83, 196]. B-spline [83, 196]. Back [69]. Back-face [69]. Background [29, 79]. Ballerinas [41]. base [135]. Based [61, 149, 151, 91, 101, 77, 221, 242, 243, 171, 66, 90, 210, 280, 140, 256, 215, 167, 85, 146, 282, 79, 272, 19, 180, 232, 13, 4, 177, 144, 276, 158]. basis [125]. Behaviour [108, 250, 23, 158]. behaviour-embedded [250]. Behavioural [90, 259, 261, 4]. between [207]. Bézier [196]. bibliography [11, 14]. Biological [50]. Biparametric [52]. Blending [103, 25]. Blends [122]. Blocks [65]. bodies [207, 24]. Body [153, 67, 161, 57, 195, 89]. Botanical [66]. boundary [49]. Bounds [139]. braids [42]. Branching [21]. Break [66]. Bridge [2]. broken [200]. broken-up [200]. bronze [281]. buffers [25]. Building [238, 239, 65]. built [278]. button [9].

Calculation [135, 154, 161]. Calder [211]. Camera [101]. Cameras [115, 254]. capability [78]. Caracol [223]. cardiac [209]. Carlo [59]. Cartesian [216]. Cartoon [129]. Casting [152, 83, 270]. Cel [129]. cell [44]. cell-transverse [44]. cellular [168, 158]. CES [105]. CG [205]. character [215]. characterization [180]. Characters [186, 260]. Choreographed [38]. Circular [103]. Classification [63]. closely [109]. Cloth [151, 205, 203]. cloth-like [203]. clothes [207]. Cloths [39].

Coherence [154, 124, 72, 93, 230, 218]. collaborative [286]. Colliding [85]. Collision [77, 69, 51, 171, 280, 24]. colour [125]. Colours [102]. Combining [85]. comment [194]. Commercial [15]. Compact [99]. company [9]. comparing [169]. Complex [50, 10, 72, 24, 249]. Complexity [104]. components [250]. Compression [141]. Computation [59, 271]. Computer [132, 129, 55, 60, 159, 156, 26, 39, 2, 137, 67, 206, 45, 41, 16, 105]. Computer-aided [156, 137]. Computer-assisted [129]. Computing [183]. Concept [3]. Concepts [57]. Conference [75]. Connectionist [12]. Connectivity [91]. Constraint [140]. Constraint-based [140]. Constraints [60, 89, 39, 267]. Constructive [232]. Contact [128, 77]. Contacts [153]. contrast [234, 235]. contrast-enhanced [234, 235]. Control [85, 206, 193, 173, 288, 257]. controlled [82]. controlling [23]. correspondences [49]. Cost [132]. Creating [62]. Creation [190, 240, 241]. Creative [35]. credibility [278]. critical [283]. Cross [74, 2]. Cross-Bridge [2]. Cross-dissolves [74]. CSG [152]. CT [242, 243]. cubic [21]. Culling [69, 218, 283]. cultural [277]. Currents [90]. Curve [99]. Curved [73]. Curves [76, 103]. cutting [236, 237]. Cylinder [127]. Cylinder-like [127]. Cylinders [76].

D [189, 222, 131, 44, 120, 217, 151, 168, 219, 78, 214, 188, 114, 47, 164, 40, 10, 272, 190, 191, 158, 16]. D-space [78]. Dangling [39]. Data [110, 53, 114, 63, 83, 107, 30, 36, 44, 40, 166]. database [47, 284]. databases [189]. decomposition [265]. Deformable [77, 209, 24, 195, 22]. Deformation [61, 97]. deformations [207]. Deformed [74, 270]. deformed-volume [270]. Density [64]. dependent [145, 114]. Derivative [99]. Design [64, 80, 26]. Desque [164]. Detail [186]. Detection [69, 51, 242, 243, 171, 280, 32]. deterministic [174]. developed [9]. Development [287]. diagnosis [209]. difference [281]. Differential [99]. Digital [112, 193]. dilemmas [278]. dimensional [68, 43, 177, 18]. Direct [70]. directed [288, 65]. disorders [286]. disparity [256]. disparity-based [256]. Display [5, 125, 44, 282, 16]. Displays [36]. dissolves [74]. Dormancy [66]. double [278]. down [157]. Driven [73, 210, 119]. droplets [157]. Dust [172]. Dynamic [133, 92, 73, 234, 235, 161, 178, 276]. Dynamics [67, 12, 34, 27].

Ecosystem [50]. Editorial [197, 264, 285, 148, 155, 160, 165, 176, 192, 208, 213, 220, 248, 252, 269, 273, 279, 96, 100, 113, 121, 117, 126, 130, 134, 138, 143, 170, 182, 258, 226, 106, 187, 233]. Efficient [152, 163, 173, 218, 180, 204]. elasticity [249]. electrostatic [219]. element [240, 241]. elements [164]. embedded [250]. embodied [262]. enabled [217]. encoded [146]. end [9]. endpoint [267]. engine [178]. enhanced [234, 235]. Environment [147, 7, 259, 288]. Environments [150, 10, 133, 72, 263, 212, 286]. Ergonomic [137]. Erosion [149]. Error [139, 228]. establishing [49]. estimation [163]. Euclidean [254]. Eurhythmy [3]. Evaluating [288]. Evaluation [137]. Evidence [186]. evolutionary [271]. Evolving [193]. experiences [278]. Exploitation [93]. Exploiting [230]. Exploration [105, 107, 33]. expressions [262]. Extended [115]. Extending [31]. Extension [154, 114, 173]. extrapolation [249]. Extrusion [118]. Eye [261].

face [49, 69]. Faces [30, 225]. Facial [31, 92, 210, 262, 13]. factorization [254]. Fast [136, 125, 280, 265, 194]. Ferret [109]. Fidelity [112, 240, 241]. Fields [149, 70, 219, 18]. Figure [88]. figures [172]. Film [101]. Filtering [53]. finite [240, 241]. Fire [68]. First [99]. Flames [68]. flexible [190]. Floor [112]. Flow [83, 38, 221]. flows [185]. focus [244, 245]. foot [177]. force [249, 177]. Forest [131]. form [82, 270, 255, 204]. Formfactor [135]. Formula [99]. Fourier [53, 141]. Fractal [94]. Frame [154, 212]. Frame-to-Frame [154]. Frames [152]. framework [184]. Free [82, 270, 255, 204]. Free-form [82, 270, 255, 204]. Freeform [122]. Full [93]. Function [221, 93]. Function-based [221]. functioning [286]. Functions [64, 10, 125, 20]. Fungal [86]. fusion [260].

G [5]. G-Octree [5]. gait [271, 287].

garments [224]. gaseous [202]. Gaze [263]. Generalized [76, 236, 237, 16]. generate [200]. generated [41]. generating [256]. Generation [53, 79, 21, 169]. generator [43, 199]. generic [178]. Genetic [88]. Geometric [132, 137]. GIS [189]. Global [110, 59]. Goal [65, 60, 288]. Goal-directed [65, 288]. goals [78]. Grammar [101]. Graphical [36, 164, 250]. Graphics [75, 105, 132, 121, 142, 9]. grasp [215]. grasp-based [215]. Grey [5]. Grey-Scale [5]. grid [44]. Growth [86, 231].

hair [25, 34]. Hanging [39]. Heating [22]. Heliotropism [66]. Hemisphere [135]. Hemisphere-base [135]. Heritage [284, 277, 275]. Hermite [21, 103]. Hertz [77]. Hierarchical [139, 230]. High [229, 112, 162, 40, 9, 286, 240, 241]. high-end [9]. high-fidelity [240, 241]. high-functioning [286]. High-quality [229, 162]. high-speed [40]. hopper [173]. hopping [172]. hull [276]. Human [271, 128, 177, 36, 209, 205, 78, 173, 161, 159, 199, 34, 231]. hundred [45]. Hyperpatch [83]. hypervolume [232].

IBCD [280]. Illumination [10, 59, 194, 144, 54]. Illusion [62]. Image [6, 111, 97, 282, 74, 38, 144, 171, 277, 280, 188, 244, 245]. Image-based [144, 171]. image-processing [277]. Images [111, 5, 242, 243, 234, 235, 281, 253, 169]. imaging [275, 181]. immersive [289, 282]. Imperfection [6]. implementing [236, 237]. implicit [180]. impostor [283]. improved [196]. Improving [249]. incorporation [283]. Incremental [145]. indexed [45]. Infinite [37]. Inflation [118]. inherent [234, 235]. ink [158]. Integrated [48, 87, 109, 40]. Integrating [207]. Interaction [120]. Interactions [56, 50, 249]. Interactive [153, 80, 203, 18, 236, 237, 227, 246, 247, 47, 284, 286, 184, 16]. interface [164, 287]. Interference [32]. Interpolating [123]. Interpolation [103, 74, 83, 175, 21, 49]. interpreting [40]. Intrinsic [123]. invariant [234, 235]. Inverse [27]. isosurfaces [162, 224]. isotropic [174]. Issue [121, 142]. Italy [8].

Jaw [128].

kidney [240, 241]. Kinematically [73]. knitwear [204]. knots [42]. knowledge [242, 243]. knowledge-based [242, 243].

L [86]. L-Systems [86]. Lagrangian [27]. Land [56]. Language [55, 63, 250]. Large [81]. Layered [257]. Leaf [102]. left [209]. length [146]. Light [95]. Lighting [212]. like [127, 203, 158]. Limb [71]. Linear [83, 175, 249, 107]. Lip [29]. Lip-Sync [29]. local [272]. locomotion [287]. LOD [257, 283]. lumigraph [257]. Lung [242, 243].

magnetic [18]. Mammal [108]. MAMVIS [108]. Manipulation [153]. Mankind [73]. mapping [266]. maps [202]. Marine [110, 108]. Mass [151]. match [205]. measurement [272]. Mechanical [58]. medical [244, 245, 181]. Medium [59]. melting [22]. Menv [7]. merging [205]. mesh [236, 237, 272]. mesh-cutting [236, 237]. Meshes [118, 227]. Metamorphosis [127]. Metaphor [101]. metastasis [242, 243]. Method [124, 150, 118, 242, 243, 200, 236, 237, 125, 254, 166, 212]. Methods [57]. Metrics [139, 169]. Minimization [132]. mirrors [281]. Mobility [218]. Mock [62]. modalities [244, 245]. Model [61, 90, 149, 151, 92, 209, 271, 173, 214, 195, 250, 240, 241]. Model-Based [151]. Modeling [202, 221, 217, 267, 225, 232]. modeller [15]. Modelling [120, 112, 76, 7, 81, 128, 95, 86, 30, 266, 24, 179, 253, 12, 34, 13, 251, 255]. Models [132, 91, 37, 123, 77, 118, 172, 282, 22, 27, 190, 251]. modification [282]. Modular [104]. molecules [219]. Monte [59]. Mood [79]. Morphing [97, 114, 256, 276]. Motion [209, 84, 140, 88, 87, 71, 128, 101, 141, 2, 65, 198, 267, 78, 215, 199, 260]. Motions [116, 79]. Motor [23]. Mountain [149]. Movement [71, 288]. movements [261]. movie [205]. movies [256]. Moving [154, 205, 157]. MR [234, 235]. Multi [57, 217, 227, 266, 212]. Multi-body [57]. Multi-Frame [212]. multi-mapping [266]. multi-resolution [217, 227]. Multidimensional [10]. Multilayer [224]. Multilevel [72]. multimedia [290, 47, 284]. Multiple [70, 186, 260, 229, 40]. Multiresolution [145]. Multivariate [63]. Muscle [2, 210]. muscle-based [210]. Museum [47, 275]. Music [79].

Natural [55, 266]. navigation [244, 245, 47]. neuromusculoskeletal [271]. Non [175, 107, 174]. non-deterministic [174]. Non-linear [175, 107]. Normal [53, 163]. novel [256, 251]. NPSNET [51]. NRC [275]. number [43].

OBB [280]. Object [120, 84, 167, 57, 21, 256, 283]. Object-order [167]. Object-oriented [120, 84, 57]. Objects [136, 154, 151, 85, 127, 144, 270, 203, 265, 178, 194]. Obstacle [153, 56]. Ocean [107, 179]. Oceanographic [111]. Octree [91, 5]. Offset [76]. Online [198, 288]. operation [201]. operator [163]. optimality [272]. Optimized [71]. Order [82, 167]. Order-controlled [82]. organizing [189]. organs [231]. Orientations [103]. oriented [277, 120, 84, 57]. Origami [98]. origin [54]. orthopaedic [244, 245].

Pacific [75, 121, 142]. Painting [129, 217]. panoramas [256]. paradigms [246, 247]. Parameter [35]. Parameters [123, 43]. Parametric [122, 214]. Parkinson [290]. Participating [59]. Particle [90, 202]. Particle-based [90]. particles [222]. Partitioning [150]. past [278]. PASTIS [17]. Path [73]. patients [290]. Patterns [151]. people [286]. Perception [19]. Perception-based [19]. Performance [210, 263, 193]. Performance-driven [210]. perfusion [234, 235]. Periodic [141]. personal [9, 41]. perspective [228, 254]. phenomena [202]. Physical [56, 290]. Physically [85, 13]. Physically-based [85, 13]. physics [224]. pixel [25]. planar [173]. Planner [71]. planning [78, 215, 23]. Playing [98]. Polygon [118]. Polyhedra [69]. Polyhedral [123]. polymer [46]. postural [78]. Posture [137]. pottery [274]. Predict [137]. presentation [284]. Priority [119]. Priority-driven [119]. Process [3, 48]. Processing [111, 186, 104, 277, 24]. processor [16]. product [9, 216]. Production [129]. program [17]. Programming [88, 58, 259]. progressive [251]. project [286, 223]. Projected [228]. Projection [93, 228, 282]. Properties [95]. Providing [110]. pseudo [282]. pseudo-projection [282]. Psychological [186]. puppetry [193].

quality [206, 229, 162]. Quaternion

[99, 103]. query [54].

R [189]. R-trees [189]. Radiative [95]. Radiosity [139, 135, 133, 72, 94, 212]. Radius [76]. Random [94]. Range [31, 281]. Rapid [225]. rasterization [32]. Ray [270, 124, 115, 152, 83, 119, 1, 54, 196]. ray-traced [1]. Reach [137]. reaction [177]. Real [186, 51, 16]. Real-time [186, 51, 16]. realism [249]. Realistic [185, 6, 188, 204, 54]. Reality [154, 62, 289, 203]. RECODE [171]. Recognition [129]. Reconstruction [112, 276, 278, 174, 229, 254, 274, 188]. rectilinear [44]. recursive [49]. references [45]. reflectances [174]. Reflection [37]. Regression [267]. regular [44]. rehabilitation [290, 289, 287]. Relativistic [136, 194]. relativity [194]. render [9]. rendered [169]. renderer [40]. Rendering [136, 131, 70, 25, 124, 114, 93, 202, 185, 179, 167, 162, 146, 218, 19, 34, 166, 196, 194, 266, 158, 283, 204]. Representation [107, 49]. Represented [5]. Reproducing [211]. research [240, 241]. resolution [217, 227]. Response [77, 51]. responsive [184]. retargetting [198]. Rigid [153, 67]. Rods [39]. Rolling [52]. run [146]. run-length-encoded [146]. running [173].

same [54]. Sampled [53]. Scalar [70]. Scale [234, 235, 5]. Scale-invariant [234, 235]. Scanned [30]. Scene [104, 276]. Scenery [131, 149, 266]. scenes [238, 239, 179, 188, 255]. Scientific [74, 11, 14]. Sculpting [191, 222]. sculpture [26]. Sea [105, 112]. seeds [33]. segmentation [246, 247, 234, 235]. Selection [35, 234, 235]. selective [166]. Sequences [111, 205, 188]. Sequential [60]. Sequential-goal [60]. set [54]. Sets [114]. Shading [63]. shadow [25]. Shape [281, 49, 272, 123, 190]. Shape-adaptive [272]. Shared [147, 263]. SharedWeb [147]. Sign [55]. silhouettes [190]. SIM [80]. Similarity [129]. Simple [158, 67]. simplification [272]. Simulating [268, 34]. Simulation [68, 66, 90, 102, 149, 56, 80, 85, 59, 2, 73, 271, 195, 158]. Simulations [58, 57]. Simulator [98, 249]. Simulators [181]. Single [255, 286]. single-user [286]. Single-view [255]. Six [45]. Six-hundred [45]. size [163]. skeletal [180]. skill [12]. Skin [92]. sky [125]. slabs [228]. Sliding [128]. Small [73, 43]. Smoke [68]. Smooth [52, 162]. snowy [266]. soft [265]. software [48]. Soil [50]. Solid [103, 15]. Solution [94]. Solutions [135]. Sound [87]. Sources [95]. Space [61, 76, 98, 280, 78, 265]. Spacetime [89]. Special [121, 142, 194]. Specification [104]. specifications [169]. specifying [23]. spectrum [286]. speech [214]. speed [40]. splashing [172]. spline [178, 83, 196]. Spread [68]. Spring [151]. Spring-Mass [151]. staging [156]. standing [288]. State [244, 245]. State-of-the-art [244, 245]. Steps [73]. stereoscopic [167]. storytelling [284]. strategy [109]. Strength [36]. String [81]. structure [34]. Subtleties [262]. Subvolume [93]. Suibokuga [158]. Suibokuga-like [158]. suited [24]. Super [80]. supervised [43]. Surface [52, 122, 168, 157, 188, 20, 180]. Surfaces [95, 94, 196]. surgery [249]. surgical [244, 245, 195]. survey [236, 237, 159]. swept [183]. Symbolic [122]. Sync [29]. Synthesis [6, 141]. System [154, 108, 21, 217, 24, 274, 289, 17, 180, 287]. Systems [57, 86, 193, 46]. task [263]. Technique [114, 205, 33, 196]. Techniques [97, 29, 236, 237, 290, 227, 219, 266]. Technology [105, 275]. Template [146, 167]. Template-based [146, 167]. temporal [218]. Terrain [112, 56, 145, 124]. TEX [80]. TEX-SIM [80]. Textile [80, 206]. Texture [81, 222, 229, 43]. Textures [131]. texturing [232]. their [168]. Theory [77]. Thickening [201]. Three [43, 177, 18]. Three-dimensional [43, 177, 18]. Tighter [139]. Time [61, 223, 114, 283, 209, 186, 51, 16]. Time-critical [283]. Time-dependent [114]. time-varying [209]. tongue [214]. tool [277, 109]. tools [236, 237]. touch [217]. touch-enabled [217]. traced [1]. Tracing [115, 119, 196]. training [181]. Trajectory [85]. transformation [49]. translating [199]. Translation [55]. Translucent [5]. transmittances [174]. transverse [44]. trauma [240, 241]. Travel [223]. Trees [66, 189, 158]. Triangulation [145]. Truth [278]. turbulent [185]. Two [68]. Two-dimensional [68]. type [116]. Types [31].

ultrasound [181]. uncalibrated [254, 188]. Unconscious [186]. Unified [39]. Unit [99]. Use [72, 135, 240, 241]. user [164, 286]. Using [222, 131, 154, 53, 97, 216, 30, 202, 236, 237, 129, 90, 125, 280, 25, 124, 81, 218, 224, 83, 163, 123, 190, 15, 287, 266].

Validity [154]. Variable [76, 163].

variable-size [163]. Various [56]. varying [209]. vascular [181]. Vector [53, 219]. Vehicles [56]. Velocity [149]. ventricle [209]. Vertical [124]. vertigo [28]. Via [110]. video [225, 199]. videoconferencing [190]. View [145, 175, 256, 276, 255]. View-dependent [145]. views [229, 256]. Virtual [66, 154, 147, 47, 98, 191, 263, 236, 237, 289, 288, 203, 28, 286, 177, 23]. virtues [28]. visibility [230, 283]. vision [206, 4]. vision-based [4]. Visual [68, 66, 90, 102, 149, 37, 55, 253, 288, 276]. Visualization [108, 91, 35, 219, 10, 46, 74, 83, 107, 242, 243, 42, 209, 48, 227, 109, 161, 281, 265, 11, 1, 14, 180, 166, 194, 18]. Visualizing [50, 20, 189]. Volume [64, 35, 70, 33, 63, 93, 270, 246, 247, 167, 40, 1, 169]. volume-rendered [169]. volumes [183, 146]. voxel [180]. voxel-based [180].

Walking [73, 159, 177]. Warping [151]. Water [90, 172, 157]. Waterfalls [81]. Web [110, 238, 239, 147]. Weighted [139]. wet [200]. while [288]. Whip [116]. Whip-type [116]. Wide [110, 147]. Within [154]. works [211]. world [177, 110, 147]. Wrinkle [92].

References

Nielson:1990:TAR

[1] Gregory M. Nielson, Thomas A. Foley, and David A. Lane. Towards animating ray-traced volume visualization. The Journal of Visualization and Computer Animation, 1(1):2–8, August 1990. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wakabayashi:1990:CSA

[2] Takeyuki Wakabayashi, Nelson Max, and Nobuhiko Hayashi. Computer simulation and animation of muscle crossbridge motion. The Journal of Visualization and Computer Animation, 1(1): 9–14, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Amkraut:1990:ECP

[3] Susan Amkraut and Michael Girard. Eurhythmy: Concept and process. The Journal of Visualization and Computer Animation, 1(1):15–17, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Thalmann:1990:VBA

[4] Daniel Thalmann, Olivier Renault, and Nadia Magnenat Thalmann. A visionbased approach to behavioural animation. The Journal of Visualization and Computer Animation, 1(1):18– 21, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kunii:1990:TDA

[5] Tosiyasu L. Kunii, Xiaoyang Mao, and Issei Fujishiro. A translucent display algorithm for G-octree represented greyscale images. The Journal of Visualization and Computer Animation, 1(1): 22–25, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Badler:1990:IRI

[6] Norman I. Badler and Welton Becket. Imperfection for realistic image synthesis. The Journal of Visualization and Computer Animation, 1(1):26– 32, August 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Leffler:1990:MMA

[7] Samuel J. Leffler, William T. Reeves, and Eben F. Ostby. The menv modelling and animation environment. The Journal of Visualization and Computer Animation, 1(1):33–40, August 1990. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

GraziaMattei:1990:AI

[8] M. Grazia Mattei. Animation in italy. The Journal of Visualization and Computer Animation, 1(2):42–43, December 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pittman:1990:RBH

[9] J. H. Pittman. The render button: how a high-end graphics company developed a personal graphics product. The Journal of Visualization and Computer Animation, 1(2):44–48, December 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Mudur:1990:MIF

[10] S. P. Mudur and Sumanta N. Pattanaik. Multidimensional illumination functions for visualization of complex 3D environments. The Journal of Visualization and Computer Animation, 1(2):49– 58, December 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Nielson:1990:ABS

[11] G. M. Nielson and K. Voegele. An annotated bibliography of scientific visualization. The Journal of Visualization and Computer Animation, 1(2):59– 65, December 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Ridsdale:1990:CMS

[12] G. Ridsdale. Connectionist modelling of skill dynamics. The Journal of Visualization and Computer Animation, 1 (2):66–72, December 1990. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Terzopoulos:1990:PBF

[13] D. Terzopoulos and K. Waters. Physicallybased facial modelling, analysis, and animation. The Journal of Visualization and Computer Animation, 1(2):73–80, December 1990. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Nielson:1991:ABS

[14] G. M. Nielson and K. Voegele. An annotated bibliography of scientific visualization. 2. The Journal of Visualization and Computer Animation, 2(1):2– 8, January–March 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wyvill:1991:CAU

[15] G. Wyvill and P. Sharp. Commercial animation using a solid modeller. The Journal of Visualization and Computer Animation, 2(1):9–15, January–March 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

vanOverveld:1991:GDP

[16] C. W. A. M. van Overveld. The generalized display processor as an approach to real-time interactive 3-D computer animation. The Journal of Visualization and Computer Animation, 2(1):16– 25, January–March 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Muller:1991:PAS

[17] H. Muller, J. Winckler, S. Grzybek, M. Otte, B. Stoll, F. Equoy, and N. Higelin. The program animation system PASTIS. The Journal of Visualization and Computer Animation, 2(1):26– 33, January–March 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yamashita:1991:IVT

[18] H. Yamashita, T. Johkoh, S. Takita, and E. Nakamae. Interactive visualization of three-dimensional magnetic fields. The Journal of Visualization and Computer Animation, 2(1):34–40, January–March 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Parke:1991:PAR

[19] F. I. Parke. Perception-based animation rendering. The Journal of Visualization and Computer Animation, 2(2): 44–51, April–June 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pottmann:1991:VFS

[20] H. Pottmann, H. Hagen, and A. Divivier. Visualizing functions on a surface. The Journal of Visualization and Computer Animation, 2(2):52–58, April–June 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

DeLeon:1991:BOG

[21] M. K. De Leon. Branching object generation and animation system with cubic Hermite interpolation. The Journal of Visualization and Computer Animation, 2(2):60–67, April–June 1991. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Terzopoulos:1991:HMD

[22] D. Terzopoulos, J. Platt, and K. Fleischer. Heating and melting deformable models. The Journal of Visualization and Computer Animation, 2(2):68–73, April–June 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zeltzer:1991:MPA

[23] David Zeltzer and Michael B. Johnson. Motor planning: an architecture for specifying and controlling the behaviour of virtual actors. The Journal of Visualization and Computer Animation, 2(2): 74–80, April–June 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). (to appear).

Gascuel:1991:MSC

[24] M.-P. Gascuel, A. Verroust, and C. Puech. A modelling system for complex deformable bodies suited to animation and collision processing. The Journal of Visualization and Computer Animation, 2(3):82–91, July–September 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

LeBlanc:1991:RHU

[25] A. M. LeBlanc, R. Turner, and D. Thalmann. Rendering hair using pixel blending and shadow buffers. The Journal of Visualization and Computer Animation, 2(3):92–97, July–September 1991. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Todd:1991:CSD

[26] S. Todd, W. Latham, and P. Hughes. Computer sculpture design and animation. The Journal of Visualization and Computer Animation, 2(3):98–105, July–September 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Vasilonikolidakis:1991:ILD

[27] N. Vasilonikolidakis and G. J. Clapworthy. Inverse Lagrangian dynamics for animating articulated models. The Journal of Visualization and Computer Animation, 2(3):106–113, July–September 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Queau:1991:VVV

[28] P. Queau. The virtues and the vertigo of the virtual. The Journal of Visualization and Computer Animation, 2 (3):114–115, July–September 1991. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Lewis:1991:ALS

[29] J. Lewis. Automated lip-sync: Background and techniques. The Journal of Visualization and Computer Animation, 2(4):118–122, October–December 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Waters:1991:MAF

[30] K. Waters and D. Terzopoulos. Modelling and animating faces using scanned data. The Journal of Visualization and Computer Animation, 2(4):123–128, October–December 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

DiPaola:1991:ERF

[31] S. DiPaola. Extending the range of facial types. The Journal of Visualization and Computer Animation, 2(4):129–131, October–December 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Shinya:1991:IDT

[32] M. Shinya and M.-C. Forgue. Interference detection through rasterization. The Journal of Visualization and Computer Animation, 2(4):132–134, October–December 1991. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kwan-LiuMa:1991:VSV

[33] Kwan-Liu Ma, M. F. Cohen, and J. S. Painter. Volume seeds: a volume exploration technique. The Journal of Visualization and Computer Animation, 2(4): 135–140, October–December 1991. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Rosenblum:1991:SSD

[34] R. E. Rosenblum, W. E. Carlson, and E. Tripp, III. Simulating the structure and dynamics of human hair: modelling, rendering and animation. The Journal of Visualization and Computer Animation, 2(4):141–148, October–December 1991. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Giertsen:1992:CPS

[35] Christopher Giertsen. Creative parameter selection for volume visualization. The Journal of Visualization and Computer Animation, 3(1):1–11, January–March 1992. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wei:1992:GDH

[36] Susanna Wei and Norman I. Badler. Graphical displays of human strength data. The Journal of Visualization and Computer Animation, 3(1):13–22, January–March 1992. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Inakage:1992:IRM

[37] Masa Inakage. Infinite reflection models for the visual arts. The Journal of Visualization and Computer Animation, 3(1):23–29, January–March 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Sims:1992:CIF

[38] Karl Sims. Choreographed image flow. The Journal of Visualization and Computer Animation, 3(1):31–43, January–March 1992. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

VanLoon:1992:HCD

[39] Erik Van Loon and C. W. A. M. Van Overveld. Hanging cloths and dangling rods: a unified approach to constraints in computer animation. The Journal of Visualization and Computer Animation, 3(1):45–59, January–March 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Miyazawa:1992:HIR

[40] T. Miyazawa and K. Koyamada. A highspeed integrated renderer for interpreting multiple 3D volume data. The Journal of Visualization and Computer Animation, 3(2):65–83, April–June 1992. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yoshimoto:1992:BGP

[41] S. Yoshimoto. Ballerinas generated by a personal computer. The Journal of Visualization and Computer Animation, 3(2):85–90, April–June 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bates:1992:VKB

[42] L. Bates, B. Wyvill, and C. Yao. The visualization of knots and braids. The Journal of Visualization and Computer Animation, 3(2):91–104, April–June 1992. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Gourret:1992:TDT

[43] J. P. Gourret and P. Afflard. Threedimensional texture generator supervised by a small number of parameters. The Journal of Visualization and Computer Animation, 3(2):105–127, April–June 1992. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Doi:1992:CDA

[44] A. Doi and A. Koide. A cell-transverse display algorithm for regular and rectilinear 3D grid data. The Journal of Visualization and Computer Animation, 3(2):129–143, April–June 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Thalmann:1992:SIR

[45] N. M. Thalmann and D. Thalmann. Six-hundred indexed references on computer animation. The Journal of Visualization and Computer Animation, 3 (3):147–174, July–September 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Munkel:1992:VPS

[46] C. Munkel and D. W. Heermann. Visualization of polymer systems. The Journal of Visualization and Computer Animation, 3(3):175–181, July–September 1992. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Miller:1992:VMI

[47] G. Miller, E. Hoffert, S. E. Chen, E. Patterson, D. Blackketter, S. Rubin, S. A. Applin, D. Yim, and J. Hannan. The virtual museum: interactive 3D navigation of a multimedia database. The Journal of Visualization and Computer Animation, 3(3):183–197, July–September 1992. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Friedell:1992:ISP

[48] M. Friedell, S. Kochhar, M. V. Lapolla, and J. Marks. Integrated software, process, algorithm and application visualization. The Journal of Visualization and Computer Animation, 3(4):201–218, October–December 1992. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Parent:1992:STB

[49] R. E. Parent. Shape transformation by boundary representation interpolation: a recursive approach to establishing face correspondences. The Journal of Visualization and Computer Animation, 3(4): 219–239, October–December 1992. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Liddell:1993:VCB

[50] C. M. Liddell and D. Hansen. Visualizing complex biological interactions in the soil ecosystem. The Journal of Visualization and Computer Animation, 4(1):3–12, January–March 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zyda:1993:NRT

[51] M. J. Zyda, D. R. Pratt, W. D. Osborne, and J. G. Monahan. NPSNET: Real-time collision detection and response. The Journal of Visualization and Computer Animation, 4(1):13–24, January–March 1993. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Hegron:1993:RSB

[52] G. Hégron. Rolling on a smooth biparametric surface. The Journal of Visualization and Computer Animation, 4 (1):25–32, January–March 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Goss:1993:NVG

[53] M. E. Goss and I. P. Page. Normal vector generation for sampled data using Fourier filtering. The Journal of Visualization and Computer Animation, 4 (1):33–50 (or 33–49??), January–March 1993. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Rappoport:1993:SOR

[54] A. Rappoport. The same origin ray set query for realistic illumination: algorithm and analysis. The Journal of Visualization and Computer Animation, 4(1):51–60, January–March 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Lee:1993:CAV

[55] J. Lee and T. L. Kunii. Computer animated visual translation from natural language to sign language. The Journal of Visualization and Computer Animation, 4(2):63–78, April–June 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Jimenez:1993:PSL

[56] S. Jimenez, A. Luciani, and O. Raoult. Physical simulation of land vehicles with obstacle avoidance and various terrain interactions. The Journal of Visualization and Computer Animation, 4(2): 79–94, April–June 1993. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kuhn:1993:AOO

[57] V. Kühn and W. Müller. Advanced object-oriented methods and concepts for simulations of multi-body systems. The Journal of Visualization and Computer Animation, 4(2):95–112 (or 95– 111??), April–June 1993. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kearney:1993:PMS

[58] J. K. Kearney, S. Hansen, and J. F. Cremer. Programming mechanical simulations. The Journal of Visualization and Computer Animation, 4(2):113–130 (or 113–129??), April–June 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pattanaik:1993:CGI

[59] Sumanta N. Pattanaik and S. P. Mudur. Computation of global illumination in a participating medium by Monte Carlo simulation. The Journal of Visualization and Computer Animation, 4(3):133–152, July–September 1993. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Liu:1993:SGC

[60] Wen-Bang Liu, Ming-Tat Ko, and Buei-Chuan Chang. Sequential-goal constraints for computer animation. The Journal of Visualization and Computer Animation, 4(3):153–164 (or 153– 163??), July–September 1993. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bechmann:1993:ATS

[61] D. Bechmann and N. Dubreuil. Animation through space and time based on a space deformation model. The Journal of Visualization and Computer Animation, 4(3):165–184, July–September 1993. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Stosser:1993:ICM

[62] A. Stösser, W. Leister, and F. Meyer. Illusion — creating a mock reality. The Journal of Visualization and Computer Animation, 4(4):187–198, October–December 1993. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Palmer:1993:LCS

[63] T. C. Palmer, E. V. Simpson, and W. M. Smith. A language for the classification and shading of multivariate volume data. The Journal of Visualization and Computer Animation, 4 (4):199–212 (or 199–211??), October– December 1993. CODEN JVCAEO.

ISSN 1049-8907 (print), 1099-1778 (electronic).

Ebert:1993:DAV

[64] D. S. Ebert. Design and animation of volume density functions. The Journal of Visualization and Computer Animation, 4(4):213–232, October–December 1993. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

vanOverveld:1993:BBG

[65] C. W. A. M. Van Overveld. Building blocks for goal-directed motion. The Journal of Visualization and Computer Animation, 4(4):233–250, October– December 1993. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Chiba:1994:VSB

[66] N. Chiba, S. Ohkawa, K. Muraoka, and M. Miura. Visual simulation of botanical trees based on virtual heliotropism and dormancy break. The Journal of Visualization and Computer Animation, 5(1):3–16 (or 3–15??), January–March 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

vanOverveld:1994:SAR

[67] C. W. A. M. Van Overveld. A simple approximation to rigid body dynamics for computer animation. The Journal of Visualization and Computer Animation, 5(1):17–36, January–March 1994. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Chiba:1994:TDV

[68] N. Chiba, K. Muraoka, H. Takahashi, and M. Miura. Two-dimensional visual simulation of flames, smoke and the spread of fire. The Journal of Visualization and Computer Animation, 5 (1):37–54 (or 37–53??), January–March 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Vanecek:1994:BFC

[69] G. Vaněček, Jr. Back-face culling applied to collision detection of polyhedra. The Journal of Visualization and Computer Animation, 5(1):55–63, January–March 1994. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Giertsen:1994:DVR

[70] C. Giertsen. Direct volume rendering of multiple scalar fields. The Journal of Visualization and Computer Animation, 5(2):69–84, April–June 1994. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Miller:1994:ALM

[71] D. P. Miller and R. E. Parent. An articulated limb motion planner for optimized movement. The Journal of Visualization and Computer Animation, 5(2):85–123, April–June 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Vilaplana:1994:MUC

[72] Josep Vilaplana and Xavier Pueyo. Multilevel use of coherence for complex radiosity environments. The Journal of Visualization and Computer Animation, 5(3):129–142 (or 129–141?), July– September 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

vanOverveld:1994:SSM

[73] C. W. A. M. van Overveld and Hyeongseok Ko. Small steps for mankind: Towards a kinematically driven dynamic simulation of curved path walking. The Journal of Visualization and Computer Animation, 5(3): 143–166, July–September 1994. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Ruprecht:1994:DCD

[74] D. Ruprecht and H. Müller. Deformed cross-dissolves for image interpolation in scientific visualization. The Journal of Visualization and Computer Animation, 5(3):167–181, July–September 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Anonymous:1994:PGC

[75] Anonymous. Pacific Graphics ’93 Conference. The Journal of Visualization and Computer Animation, 5(4):??, October–December 1994. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kim:1994:MAG

[76] Myung-Soo Kim, Eun-Joo Park, and Hwan-Yong Lee. Modelling and animation of generalized cylinders with variable radius offset space curves. The Journal of Visualization and Computer Animation, 5(4):189–207 (or 189– 208??), October–December 1994. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Terasawa:1994:CRD

[77] M. Terasawa and F. Kimura. Collision response for deformable models based on Hertz’s contact theory. The Journal of Visualization and Computer Animation, 5(4):209–224, October–December 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Jung:1994:AHA

[78] Moon-Ryul Jung, N. Badler, and T. Noma. Animated human agents with motion planning capability for 3D-space postural goals. The Journal of Visualization and Computer Animation, 5(4): 225–246, October–December 1994. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Nakamura:1994:ABM

[79] J.-I. Nakamura, T. Kaku, K. Hyun, T. Noma, and S. Yoshida. Automatic background music generation based on actors’ mood and motions. The Journal of Visualization and Computer Animation, 5(4):247–264, October–December 1994. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kambe:1994:STS

[80] H. Kambe, N. Yamasaki, and Myeong Won Lee. Super TEX-SIM: Interactive textile design and simulation. The Journal of Visualization and Computer Animation, 5(4):265–279 (or 265–280??), October–December 1994. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Mallinder:1995:MLW

[81] H. Mallinder. The modelling of large waterfalls using string texture. The Journal of Visualization and Computer Animation, 6(1):3–10, January–March 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bechmann:1995:OCF

[82] D. Bechmann and N. Dubreuil. Ordercontrolled free-form animation. The Journal of Visualization and Computer Animation, 6(1):11–32, January–March 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Semwal:1995:ARC

[83] S. K. Semwal and P. G. Swann. Applying ray casting to the flow visualization data using linear and B-spline hyperpatch interpolation. The Journal of Visualization and Computer Animation, 6(1):33–48 (or 33–47??), January–March 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Erkan:1995:OOM

[84] B. Erkan and B. Özgüç. Object-oriented motion abstraction. The Journal of Visualization and Computer Animation, 6 (1):49–66 (or 49–65??), January–March 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Lamouret:1995:CPB

[85] A. Lamouret, M.-P. Gascuel, and J.- D. Gascuel. Combining physically-based simulation of colliding objects with trajectory control. The Journal of Visualization and Computer Animation, 6(2): 71–90, April–June 1995. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Tunbridge:1995:SAM

[86] A. Tunbridge and H. Jones. An Lsystems approach to the modelling of fungal growth. The Journal of Visualization and Computer Animation, 6(2): 91–108 (or 91–107??), April–June 1995.

CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Hahn:1995:IAM

[87] J. K. Hahn, J. Geigel, Jong Won Lee, L. Gritz, T. Takala, and S. Mishra. An integrated approach to motion and sound. The Journal of Visualization and Computer Animation, 6(2):109–124 (or 109–123??), April–June 1995. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Gritz:1995:GPA

[88] L. Gritz and J. K. Hahn. Genetic programming for articulated figure motion. The Journal of Visualization and Computer Animation, 6(3):129–142, July– September 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (elec-

tronic). URL ftp://ftp.ira.uka.
de/pub/bibliography/Ai/genetic.
programming.bib.gz;   http://www.
seas.gwu.edu/student/gritz/index.
html;   http://www.seas.gwu.edu/
faculty/hahn/.

Tang:1995:BSC

[89] D. Tang, J. T. Ngo, and J. Marks. Nbody spacetime constraints. The Journal of Visualization and Computer Animation, 6(3):143–154, July–September 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Chiba:1995:VSW

[90] N. Chiba, S. Sanakanishi, K. Yokoyama, I. Ootawara, K. Muraoka, and N. Saito. Visual simulation of water currents using a particle-based behavioural model. The Journal of Visualization and Computer Animation, 6(3):155–172 (or 155– 171??), July–September 1995. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Fedenczuk:1995:VCO

[91] L. L. Fedenczuk and B. Wyvill. Visualization of connectivity in octree based models. The Journal of Visualization and Computer Animation, 6(3):173–190 (or 173–189??), July–September 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wu:1995:DWM

[92] Yin Wu, N. Magnenat Thalmann, and D. Thalmann. A dynamic wrinkle model in facial animation and skin ageing. The Journal of Visualization and Computer Animation, 6(4):195–206 (or 195– 205??), October–December 1995. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zhou:1995:SPF

[93] Yong Zhou and Zesheng Tang. Subvolume projection: A full exploitation of function coherence for volume rendering. The Journal of Visualization and Computer Animation, 6(4):207–218, October–December 1995. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Wu:1995:RSR

[94] En-Hua Wu. A radiosity solution for random fractal surfaces. The Journal of Visualization and Computer Animation, 6(4):219–230 (or 219–229??), October– December 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Paul:1995:MRP

[95] J. C. Paul, P. M. Deville, and C. Winkler. Modelling radiative properties of light sources and surfaces. The Journal of Visualization and Computer Animation, 6(4):231–240, October–December 1995. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

MagnenatThalmann:1996:Ea

[96] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 7(1):1, January 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=61006160.

Lee:1996:IMU

[97] Seung-Yong Lee, Kyung-Yong Chwa, James Hahn, and Sung Yong Shin. Image morphing using deformation techniques. The Journal of Visualization and Computer Animation, 7(1):3–24 (or 3–23??), January 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=61006161.

Miyazaki:1996:OPS

[98] Shin-Ya Miyazaki, Takami Yasuda, Shigeki Yokoi, and Jun-Ichiro Toriwaki. An origami playing simulator in the virtual space. The Journal of Visualization and Computer Animation, 7(1):25–42, January 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=61006159.

Kim:1996:CDF

[99] Myoung-Jun Kim, Myung-Soo Kim, and Sung Yong Shin. A compact differential formula for the first derivative of a unit quaternion curve. The Journal of Visualization and Computer Animation, 7(1):43–58 (or 43– 57??), January 1996. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=61006158.

MagnenatThalmann:1996:Eb

[100] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 7(2):ii, April 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23591.

Palamidese:1996:CMM

[101] Patrizia Palamidese. A camera motion metaphor based on film grammar. The Journal of Visualization and Computer Animation, 7(2):61–78, April 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23595.

Chiba:1996:VSL

[102] Norishige Chiba, Ken Ohshida, Kazunobu Muraoka, and Nobuji Saito. Visual simulation of leaf arrangement and autumn colours. The Journal of Visualization and Computer Animation, 7(2): 79–94 (or 79–93??), April 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=23594.

Kim:1996:HIS

[103] Myung-Soo S. Kim and Kee-Won W. Nam. Hermite interpolation of solid orientations with circular blending quaternion curves. The Journal of Visualization and Computer Animation, 7(2):95–110, April 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=23593.

Sun:1996:CAM

[104] Hanqiu Sun. Complexity analysis of modular specification and processing of scene animations. The Journal of Visualization and Computer Animation, 7(2):111–120 (or 111– 119??), April 1996. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=23592.

Anonymous:1996:CGT

[105] Anonymous. Computer Graphics Technology for the Exploration of the Sea CES ’95. The Journal of Visualization and Computer Animation, 7(3):??, July 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Urban:1996:E

[106] Bodo Urban. Editorial. The Journal of Visualization and Computer Animation, 7(3):123–124, July 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=23597.

Tuohy:1996:NLD

[107] S. T. Tuohy and N. M. Patrikalakis.

Non-linear data representation for ocean exploration and visualization. The Journal of Visualization and Computer Animation, 7(3):125–139 (or 135– 140??), July 1996. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23598.

Fedak:1996:MMM

[108] M. A. Fedak, P. Lovell, and B. J.

McConnell. MAMVIS: A marine mammal behaviour visualization system. The Journal of Visualization and Computer Animation, 7(3):141– 147 (or 141–148??), July 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23599.

Hankin:1996:STF

[109] S. Hankin, D. E. Harrison, J. Osborne, J. Davison, and K. O’Brien. A strategy and a tool, Ferret, for closely integrated visualization and analysis. The Journal of Visualization and Computer Animation, 7(3):149– 157 (or 149–158??), July 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23600.

Dodge:1996:PGA

[110] Chris Dodge, Frank Majewski, Beate Marx, Hans Pfeiffenberger, and Manfred Reinke. Providing global access to marine data via the World Wide Web. The Journal of Visualization and Computer Animation, 7(3):159–168, July 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23601.

Herlin:1996:IPS

[111] I. L. Herlin, I. Cohen, and S. Bouzidi. Image processing for sequences of oceanographic images. The Journal of Visualization and Computer Animation, 7(3):169–176, July 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23602.

Falcidieno:1996:HFD

[112] Bianca Falcidieno, Stefano Orgolesu, Corrado Pizzi, Antonella Sanguineti, and Michela Spagnuolo. High fidelity digital terrain modelling for the reconstruction of the Antarctic sea floor. The Journal of Visualization and Computer Animation, 7(3):177– 187 (or 177–188??), July 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23603.

MagnenatThalmann:1996:Ec

[113] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 7(4):191, October 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23608.

Lucas:1996:TDD

[114] L. Lucas, F. Trunde, and N. Bonnet.

Time-dependent 3D data sets rendering: An extension of the morphing technique. The Journal of Visualization and Computer Animation, 7(4):193–210 (or 193–209??), October 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23606.

Loffelmann:1996:RTE

[115] Helwig Löffelmann and Eduard Gröller.

Ray tracing with extended cameras. The Journal of Visualization and Computer Animation, 7(4):211–228 (or 211–227??), October 1996. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23607.

Bhat:1996:AWT

[116] Dinkar N. Bhat and Joseph K. Kearney. On animating whip-type motions. The Journal of Visualization and Computer Animation, 7(4):229–250 (or 229–249??), October 1996. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=23605.

MagnenatThalmann:1997:Ea

[117] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 8(1):1, January 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14189;   http://

www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14189&PLACEBO=IE.
pdf.

VanOverveld:1997:PIA

[118] C. W. A. M. Van Overveld and B. Wyvill. Polygon inflation for animated models: A method for the extrusion of arbitrary polygon meshes. The Journal of Visualization and Computer Animation, 8(1): 3–16, January 1997. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14188;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14188&PLACEBO=IE.
pdf.

Yagel:1997:PDR

[119] Roni Yagel and John Meeker. Prioritydriven ray tracing. The Journal of Visualization and Computer Animation, 8(1):17–32, January 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14187;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14187&PLACEBO=IE.
pdf.

Dollner:1997:OOM

[120] Jürgen Döllner and Klaus Hinrichs. Object-oriented 3D modelling, animation and interaction. The Journal of Visualization and Computer Animation, 8(1):33–64, January 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14186;   http://

www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14186&PLACEBO=IE.
pdf.

MagnenatThalmann:1997:ESI

[121] Nadia Magnenat Thalmann, Daniel Thalmann, Sung Yong Shin, Tosiyasu L. Kunii, and Myung-Soo Kim. Editorial: Special issue of Pacific Graphics 95. The Journal of Visualization and Computer Animation, 8(2): 67–68, March 1997. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14191;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14191&PLACEBO=IE.
pdf.

Kim:1997:SAF

[122] Kwansik Kim and Gershon Elber. A symbolic approach to freeform parametric surface blends. The Journal of Visualization and Computer Animation, 8(2):69–80, March 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14192;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14192&PLACEBO=IE.
pdf.

Sun:1997:IPM

[123] Yue Man Sun, Wenping Wang, and Francis Y. L. Chin. Interpolating polyhedral models using intrinsic shape parameters. The Journal of Visualization and Computer Animation, 8(2):81–96, March 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14193;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14193&PLACEBO=IE.
                                        pdf.

Lee:1997:TRM

[124] Cheol-Hi Lee and Yeong Gil Shin. A terrain rendering method using vertical ray coherence. The Journal of Visualization and Computer Animation, 8(2): 97–114 (or 97–112??), March 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14194;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14194&PLACEBO=IE.
                                        pdf.

Dobashi:1997:FDM

[125] Yoshinori Dobashi, Tomoyuki Nishita, Kazufumi Kaneda, and Hideo Yamashita. A fast display method of sky colour using basis functions. The Journal of Visualization and Computer Animation, 8(2):115–127, March 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14195;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14195&PLACEBO=IE.
                                        pdf.

MagnenatThalmann:1997:Eb

[126] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 8(3):129, July 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-

bin/abstract?ID=14200;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14200&PLACEBO=IE.
pdf.

Lazarus:1997:MCL

[127] Francis Lazarus and Anne Verroust.

Metamorphosis of cylinder-like objects. The Journal of Visualization and Computer Animation, 8(3): 131–146, July 1997. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14197;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14197&PLACEBO=IE.
pdf.

Myszkowski:1997:MHJ

[128] Karol Myszkowski, Oleg G. Okunev, Tosiyasu L. Kunii, and Masumi Ibusuki. Modelling of human jaw motion in sliding contact. The Journal of Visualization and Computer Animation, 8(3):147–163, July 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14199;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14199&PLACEBO=IE.
pdf.

Chang:1997:ACP

[129] Chueh-Wei Chang and Suh-Yin Lee.

Automatic cel painting in computerassisted cartoon production using similarity recognition. The Journal of Visualization and Computer Animation, 8(3):165–185, July 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14198;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14198&PLACEBO=IE.
                                        pdf.

MagnenatThalmann:1997:Ec

[130] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 8(4):189, October–December 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14205;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14205&PLACEBO=IE.
                                        pdf.

Chiba:1997:RFS

[131] Norishige Chiba, Kazunobu Muraoka, Akio Doi, and Junya Hosokawa. Rendering of forest scenery using 3D textures. The Journal of Visualization and Computer Animation, 8(4):191– 199, October–December 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=14202;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=14202&PLACEBO=IE.
                                        pdf.

Breen:1997:CMA

[132] David E. Breen. Cost minimization for animated geometric models in computer graphics. The Journal of Visualization and Computer Animation, 8(4): 201–220, October–December 1997. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-

bin/abstract?ID=14204;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14204&PLACEBO=IE.
pdf.

Pueyo:1997:RDE

[133] Xavier Pueyo, Dani Tost, Ignacio Martin, and Blanca Garcia. Radiosity for dynamic environments. The Journal of Visualization and Computer Animation, 8(4):221–231, October–December 1997. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=14203;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=14203&PLACEBO=IE.
pdf.

MagnenatThalmann:1998:Ea

[134] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 9(1):1, January–March 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=5504;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=5504&PLACEBO=IE.
pdf.

Doi:1998:ARS

[135] Akio Doi and Takayuki Itoh. Acceleration radiosity solutions through the use of hemisphere-base formfactor calculation. The Journal of Visualization and Computer Animation, 9(1): 3–15, January–March 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-

                                        bin/abstract?ID=5501;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=5501&PLACEBO=IE.
                                        pdf.

Betts:1998:FRR

[136] Christopher Betts. Fast rendering of relativistic objects. The Journal of Visualization and Computer Animation, 9 (1):17–31, January–March 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=5502;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=5502&PLACEBO=IE.
                                        pdf.

Wang:1998:GAP

[137] Xuguang Wang and Jean Pierre Verriest. A geometric algorithm to predict the arm reach posture for computer-aided ergonomic evaluation. The Journal of Visualization and Computer Animation, 9(1):33–47, January–March 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=5503;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=5503&PLACEBO=IE.
                                        pdf.

MagnenatThalmann:1998:Eb

[138] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 9(2):51, April–June 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=5509;   http://
                                        www3.interscience.wiley.com/cgi-

bin/fulltext?ID=5509&PLACEBO=IE.
pdf.

Chang:1998:TEB

[139] Chin-Chen Chang and Zen-Chung Shih.

Tighter error bounds and weighted error metrics for hierarchical radiosity. The Journal of Visualization and Computer Animation, 9(2):53– 63, April–June 1998. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=5506;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=5506&PLACEBO=IE.
pdf.

Gleicher:1998:CBM

[140] Michael Gleicher and Peter Litwinowicz. Constraint-based motion adaptation. The Journal of Visualization and Computer Animation, 9(2): 65–94, April–June 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=5507;   http://
www3.interscience.wiley.com/cgi-
bin/fulltext?ID=5507&PLACEBO=IE.
pdf.

Shinya:1998:PMS

[141] Mikio Shinya, Takeaki Mori, and Noriyoshi Osumi. Periodic motion synthesis and Fourier compression. The Journal of Visualization and Computer Animation, 9(2): 95–107, April–June 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

/www3.interscience.wiley.com/cgi-
bin/abstract?ID=5508;   http://

                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=5508&PLACEBO=IE.
                                        pdf.

Thalmann:1998:SIP

[142] N. M. Thalmann and D. Thalmann. Special issue on Pacific Graphics ’97. The Journal of Visualization and Computer Animation, 9(3):111–??, July– September 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

MagnenatThalmann:1998:Ec

[143] Nadia Magnenat Thalmann, Daniel Thalmann, Yeong Gil Shin, and James K. Hahn. Editorial. The Journal of Visualization and Computer Animation, 9(3):111–112, July– September 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract?ID=10005682;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=10005682&PLACEBO=
                                        IE.pdf.

Wong:1998:IIB

[144] Tien-Tsin Wong, Pheng-Ann Heng, Siu-Hang Or, and Wai-Yin Ng. Illumination of image-based objects. The Journal of Visualization and Computer Animation, 9(3):113–127, July–September 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10005681;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=10005681&
                                        PLACEBO=IE.pdf.

Klein:1998:IVD

[145] Reinhard Klein, Daniel Cohen-Or, and Tobias Hüttner. Incremental viewdependent multiresolution triangulation of terrain. The Journal of Visualization and Computer Animation, 9(3):129–143, July–September 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=10005678;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=10005678&
PLACEBO=IE.pdf.

Lee:1998:TBR

[146] Cheol-Hi Lee, Yun-Mo Koo, and Yeong Gil Shin. Template-based rendering of run-length-encoded volumes. The Journal of Visualization and Computer Animation, 9(3): 145–161, July–September 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=10005679;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=10005679&
PLACEBO=IE.pdf.

Huang:1998:SSV

[147] Jiung-Yao Huang, Chao-Tsou Fang-Tsou, Jia-Lin Chang, and Ai-Jye Lee. SharedWeb — A shared virtual environment over the World Wide Web. The Journal of Visualization and Computer Animation, 9(3): 163–182, July–September 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=10005680;

                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=10005680&
                                        PLACEBO=IE.pdf.

Magnenat-Thalmann:1998:E

[148] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 9(4):183, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=30002541;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=30002541&
                                        PLACEBO=IE.pdf.

Chiba:1998:EMB

[149] N. Chiba, K. Muraoka, and K. Fujita. An erosion model based on velocity fields for the visual simulation of mountain scenery. The Journal of Visualization and Computer Animation, 9 (4):185–194, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=30002535;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=30002535&
                                        PLACEBO=IE.pdf.

Meneveaux:1998:NPM

[150] D. Meneveaux, K. Bouatouch, E. Maisel, and R. Delmont. A new partitioning method for architectural environments. The Journal of Visualization and Computer Animation, 9(4):195– 213, October–December 1998. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=30002536;

 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=30002536&
PLACEBO=IE.pdf.

Fan:1998:SMM

[151] Jin Fan, Qifu Wang, Shiang-Fong Chen, Matthew M. F. Yuen, and C. C. Chan. A spring-mass model-based approach for warping cloth patterns on 3D objects. The Journal of Visualization and Computer Animation, 9 (4):215–227, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=30002537;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=30002537&
PLACEBO=IE.pdf.

Sanna:1998:EAR

[152] Andrea Sanna and Paolo Montuschi. An efficient algorithm for ray casting of CSG animation frames. The Journal of Visualization and Computer Animation, 9 (4):229–242, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=30002538;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=30002538&
PLACEBO=IE.pdf.

Buck:1998:IRB

[153] Matthias Buck and Elmar Schömer. Interactive rigid body manipulation with obstacle contacts. The Journal of Visualization and Computer Animation, 9 (4):243–257, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.

                                        com/cgi-bin/abstract?ID=30002539;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=30002539&
                                        PLACEBO=IE.pdf.

Duckstein:1998:EVC

[154] F. Duckstein. Extension of validity calculation to moving objects within a virtual reality system using frame-to-frame coherence. The Journal of Visualization and Computer Animation, 9 (4):259–272, October–December 1998. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=30002540;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=30002540&
                                        PLACEBO=IE.pdf.

Magnenat-Thalmann:1999:Ea

[155] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 10(1):1, January–March 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=60501347;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=60501347&
                                        PLACEBO=IE.pdf.

Palamidese:1999:CAS

[156] Patrizia Palamidese. Computer-aided staging. The Journal of Visualization and Computer Animation, 10 (1):3–14, January–March 1999. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=60501343;
                                         http://www3.interscience.wiley.

com/cgi-bin/fulltext?ID=60501343&
PLACEBO=IE.pdf.

Kaneda:1999:AWD

[157] Kazufumi Kaneda, Shinya Ikeda, and Hideo Yamashita. Animation of water droplets moving down a surface. The Journal of Visualization and Computer Animation, 10(1): 15–26, January–March 1999. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=60501344;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=60501344&
PLACEBO=IE.pdf.

Zhang:1999:SCA

[158] Qing Zhang, Youetsu Sato, Jun ya Takahashi, Kazunobu Muraoka, and Norishige Chiba. Simple cellular automatonbased simulation of ink behaviour and its application to Suibokuga-like 3D rendering of trees. The Journal of Visualization and Computer Animation, 10(1):27–37, January–March 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=60501345;
 http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=60501345&
PLACEBO=IE.pdf.

Multon:1999:CAH

[159] Franck Multon, Laure France, Marie-Paule Cani-Gascuel, and Giles Debunne. Computer animation of human walking: a survey. The Journal of Visualization and Computer Animation, 10(1):39–54, January–March 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=60501346;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=60501346&
                                        PLACEBO=IE.pdf.

Magnenat-Thalmann:1999:Eb

[160] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 10(2):55, April–June 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=62001074;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=62001074&
                                        PLACEBO=IE.pdf.

Komura:1999:CVD

[161] Taku Komura, Yoshihisa Shinagawa, and Tosiyasu L. Kunii. Calculation and visualization of the dynamic ability of the human body. The Journal of Visualization and Computer Animation, 10(2):57–78, April–June 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

                                        http://www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=62001070;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=62001070&
                                        PLACEBO=IE.pdf.

LaMar:1999:HQR

[162] Eric LaMar, Bernd Hamann, and Kenneth I. Joy. High-quality rendering of smooth isosurfaces. The Journal of Visualization and Computer Animation, 10(2):79–90, April–June 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=62001071;
http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=62001071&
PLACEBO=IE.pdf.

Shin:1999:ENE

[163] Byeong-Seok Shin. Efficient normal estimation using variable-size operator. The Journal of Visualization and Computer Animation, 10 (2):91–107, April–June 1999. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=62001073;
http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=62001073&
PLACEBO=IE.pdf.

Miller:1999:IEG

[164] Gavin Miller, Sally Grisedale, and Kenneth T. Anderson. 3Desque: interface elements for a 3D graphical user interface. The Journal of Visualization and Computer Animation, 10(2):109–119, April–June 1999. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL

http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=62001072;
http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=62001072&
PLACEBO=IE.pdf.

Magnenat-Thalmann:1999:Ec

[165] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 10(3):121, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/

abstract/67501520/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=67501520&PLACEBO=
IE.pdf.

Wang:1999:SRM

[166] Wencheng Wang, Enhua Wu, and Nelson Max. A selective rendering method for data visualization. The Journal of Visualization and Computer Animation, 10(3):123–131, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/67501521/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=67501521&PLACEBO=
IE.pdf.

Koo:1999:OOT

[167] Yun-Mo Koo, Cheol-Hi Lee, and Yeong-Gil Shin. Object-order templatebased approach for stereoscopic volume rendering. The Journal of Visualization and Computer Animation, 10(3):133–142, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/67501522/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=67501522&PLACEBO=
IE.pdf.

Gobron:1999:SCA

[168] Stéphane Gobron and Norishige Chiba. 3D surface cellular automata and their applications. The Journal of Visualization and Computer Animation, 10(3):143–158, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/

abstract/67501523/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=67501523&PLACEBO=
IE.pdf.

Williams:1999:MGS

[169] Peter L. Williams and Samuel P. Uselton. Metrics and generation specifications for comparing volume-rendered images. The Journal of Visualization and Computer Animation, 10 (3):159–178, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/67501524/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=67501524&PLACEBO=
IE.pdf.

MagnenatThalmann:1999:E

[170] Nadia Magnenat Thalmann, Yoshihisa Shinagawa, and Hung Chuan Teh. Editorial. The Journal of Visualization and Computer Animation, 10(4):179–180, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/68501004/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=68501004&PLACEBO=
IE.pdf.

Baciu:1999:RIB

[171] George Baciu, Wingo Sai-Keung Wong, and Hanqiu Sun. RECODE: an image-based collision detection algorithm. The Journal of Visualization and Computer Animation, 10 (4):181–192, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/68501005/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=68501005&PLACEBO=
                                        IE.pdf.

Ashraf:1999:DWS

[172] Golam Ashraf and Kok Cheong Wong. Dust and water splashing models for hopping figures. The Journal of Visualization and Computer Animation, 10(4):193–213, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/68501006/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=68501006&PLACEBO=
                                        IE.pdf.

Kang:1999:ECH

[173] Young-Min Kang, Hwan-Gue Cho, and Ee-Taek Lee. An efficient control over human running animation with extension of planar hopper model. The Journal of Visualization and Computer Animation, 10(4):215–224, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/68501003/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=68501003&PLACEBO=
                                        IE.pdf.

Baranoski:1999:NDR

[174] Gladimir V. G. Baranoski and Jon G. Rokne. A non-deterministic reconstruction approach for isotropic reflectances and transmittances. The Journal of Visualization and Computer Animation, 10(4):225–231, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/68501007/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=68501007&PLACEBO=
IE.pdf.

Bao:1999:NLV

[175] Hujun Bao, Li Chen, Jianguo Ying, and Qunsheng Peng. Non-linear view interpolation. The Journal of Visualization and Computer Animation, 10(4):233–241, 1999. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/68501008/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=68501008&PLACEBO=
IE.pdf.

Magnenat-Thalmann:2000:Ea

[176] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 11(1):1, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/71004331/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=71004331&PLACEBO=
IE.pdf.

Tsutsuguchi:2000:HWA

[177] Ken Tsutsuguchi, Satoshi Shimada, Yasuhito Suenaga, Noboru Sonehara, and Sakuichi Ohtsuka. Human walking animation based on foot reaction force in the three-dimensional virtual world. The Journal of Visualization and Computer Animation, 11 (1):3–16, 2000. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/71004330/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=71004330&PLACEBO=
                                        IE.pdf.

Remion:2000:DAE

[178] Yannick Rémion, Jean-Michel Nourrit, and Didier Gillard. A dynamic animation engine for generic spline objects. The Journal of Visualization and Computer Animation, 11(1):17–26, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/71004333/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=71004333&PLACEBO=
                                        IE.pdf.

Gonzato:2000:MRO

[179] Jean-Christophe Gonzato and Bertrand Le Saëc. On modelling and rendering ocean scenes. The Journal of Visualization and Computer Animation, 11(1):27–37, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/71004332/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=71004332&PLACEBO=
                                        IE.pdf.

Prevost:2000:EVB

[180] Stéphanie Prévost and Laurent Lucas. An efficient voxel-based visualization system from an implicit skeletal surface characterization. The Journal of Visualization and Computer Animation, 11(1):39–49, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/71004328/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=71004328&PLACEBO=
IE.pdf.

Troccaz:2000:SMT

[181] Jocelyne Troccaz, Delphine Henry, Noureddine Laieb, Guillaume Champleboux, Jean-Luc Bosson, and Olivier Pichot. Simulators for medical training: application to vascular ultrasound imaging. The Journal of Visualization and Computer Animation, 11(1):51–65, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/71004329/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=71004329&PLACEBO=
IE.pdf.

MagnenatThalmann:2000:E

[182] Nadia Magnenat Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 11(2):67, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72504118/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72504118&PLACEBO=
IE.pdf.

Abrams:2000:CSV

[183] Steven Abrams and Peter K. Allen.

Computing swept volumes. The Journal of Visualization and Computer Animation, 11(2):69–82, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/72504114/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=72504114&PLACEBO=
                                        IE.pdf.

Sun:2000:FIR

[184] Hanqiu Sun and Mark Green. A framework for interactive responsive animation. The Journal of Visualization and Computer Animation, 11(2):83–94, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/72504115/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=72504115&PLACEBO=
                                        IE.pdf.

Arques:2000:RRT

[185] Didier Arquès, Eric Felgines, Sylvain Michelin, and Karine Zampieri. Realistic rendering with turbulent flows. The Journal of Visualization and Computer Animation, 11(2): 95–104, 2000. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/72504116/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=72504116&PLACEBO=
                                        IE.pdf.

Oesker:2000:PEU

[186] Markus Oesker, Heiko Hecht, and Bernhard Jung. Psychological evidence for unconscious processing of detail in real-time animation of multiple characters. The Journal of Visualization and Computer Animation, 11(2):105–112, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72504117/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72504117&PLACEBO=
IE.pdf.

Wolter:2000:E

[187] Franz-Erich Wolter and Nicholas M.

Patrikalakis. Editorial. The Journal of Visualization and Computer Animation, 11(3):113, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72515295/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72515295&PLACEBO=
IE.pdf.

Koch:2000:RSR

[188] Reinhard Koch, Marc Pollefeys, and Luc Van Gool. Realistic surface reconstruction of 3D scenes from uncalibrated image sequences. The Journal of Visualization and Computer Animation, 11(3):115–127, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72515296/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72515296&PLACEBO=
IE.pdf.

Kofler:2000:RTO

[189] Michael Kofler, Michael Gervautz, and Michael Gruber. R-trees for organizing and visualizing 3D GIS databases. The Journal of Visualization and Computer Animation, 11(3):129–143, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72515299/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72515299&PLACEBO=
IE.pdf.

Weik:2000:CFA

[190] S. Weik, J. Wingbermühle, and W. Niem. Creation of flexible anthropomorphic models for 3D videoconferencing using shape from silhouettes. The Journal of Visualization and Computer Animation, 11 (3):145–154, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72515298/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72515298&PLACEBO=
IE.pdf.

Wong:2000:VS

[191] Janis P. Y. Wong, Rynson W. H. Lau, and Lizhuang Ma. Virtual 3D sculpting. The Journal of Visualization and Computer Animation, 11(3):155–166, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/72515297/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=72515297&PLACEBO=
IE.pdf.

Magnenat-Thalmann:2000:Eb

[192] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 11(4):167, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/73502729/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=73502729&PLACEBO=
IE.pdf.

Gildfind:2000:EPC

[193] Andrew Gildfind, Michael A. Gigante, and Ghassan Al-Qaimari. Evolving performance control systems for digital puppetry. The Journal of Visualization and Computer Animation, 11(4):169–183, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/73502730/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=73502730&PLACEBO=
IE.pdf.

Weiskopf:2000:IAV

[194] Daniel Weiskopf, Ute Kraus, and Hanns Ruder. Illumination and acceleration in the visualization of special relativity: a comment on fast rendering of relativistic objects. The Journal of Visualization and Computer Animation, 11(4):185–195, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/73502727/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=73502727&PLACEBO=
IE.pdf.

Meseure:2000:DBM

[195] P. Meseure and C. Chaillou. A deformable body model for surgical simulation. The Journal of Visualization and Computer Animation, 11(4):197–208, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/73502728/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=73502728&PLACEBO=
                                        IE.pdf.

Wang:2000:IRT

[196] Shyue-Wu Wang, Zen-Chung Shih, and Ruei-Chuan Chang. An improved rendering technique for ray tracing Bézier and B-spline surfaces. The Journal of Visualization and Computer Animation, 11(4):209–219, 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/73502731/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=73502731&PLACEBO=
                                        IE.pdf.

Anonymous:2000:E

[197] Anonymous. Editorial. The Journal of Visualization and Computer Animation, 11(5):221–222, ???? 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/76506848/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=76506848&PLACEBO=
                                        IE.pdf.

Choi:2000:OMR

[198] Kwang-Jin Choi and Hyeong-Seok Ko. Online motion retargetting. The Journal of Visualization and Computer Animation, 11(5):223–235, ???? 2000. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/76506845/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=76506845&PLACEBO=
IE.pdf.

Noma:2000:MGA

[199] Tsukasa Noma, Kyoji Oishi, Hiroshi Futsuhara, Hiromi Baba, Takeshi Ohashi, and Toshiaki Ejima. A motion generator approach to translating human motion from video to animation. The Journal of Visualization and Computer Animation, 11(5): 237–248, ???? 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/76506844/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=76506844&PLACEBO=
IE.pdf.

Bruderlin:2000:MGW

[200] Armin Bruderlin. A method to generate wet and broken-up animal fur. The Journal of Visualization and Computer Animation, 11(5):249– 259, ???? 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/76506849/START;   http:/
/www3.interscience.wiley.com/cgi-

                                        bin/fulltext?ID=76506849&PLACEBO=
                                        IE.pdf.

Brandel:2000:TOA

[201] Sylvain Brandel, Dominique Bechmann, and Yves Bertrand. Thickening: an operation for animation. The Journal of Visualization and Computer Animation, 11(5):261–277, ???? 2000. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/76506846/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=76506846&PLACEBO=
                                        IE.pdf.

Adabala:2000:MRG

[202] Neeharika Adabala and Swami Manohar. Modeling and rendering of gaseous phenomena using particle maps. The Journal of Visualization and Computer Animation, 11(5):279–293, ???? 2000. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http:/

                                        /www3.interscience.wiley.com/cgi-
                                        bin/abstract/76506850/START;   http:
                                        //www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=76506850&PLACEBO=
                                        IE.pdf.

Meyer:2001:IAC

[203] Mark Meyer, Gilles Debunne, Mathieu Desbrun, and Alan H. Barr. Interactive animation of cloth-like objects in virtual reality. The Journal of Visualization and Computer Animation, 12 (1):1–12, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/82003014/START;   http:/
                                        /www3.interscience.wiley.com/cgi-

bin/fulltext?ID=82003014&PLACEBO=
IE.pdf.

Zhong:2001:RER

[204] Hua Zhong, Ying-Qing Xu, Baining Guo, and Heung-Yeung Shum. Realistic and efficient rendering of freeform knitwear. The Journal of Visualization and Computer Animation, 12 (1):13–22, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/82003015/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=82003015&PLACEBO=
IE.pdf.

Hoshino:2001:MMT

[205] Jun’ichi Hoshino, Masanobu Yamamoto, and Hirofumi Saito. A match moving technique for merging CG cloth and human movie sequences. The Journal of Visualization and Computer Animation, 12(1):23–29, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/82003011/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=82003011&PLACEBO=
IE.pdf.

Anagnostopoulos:2001:CVA

[206] C. Anagnostopoulos, D. Vergados, E. Kayafas, V. Loumos, and G. Stassinopoulos. A computer vision approach for textile quality control. The Journal of Visualization and Computer Animation, 12(1):31–44, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/

                                        abstract/82003016/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=82003016&PLACEBO=
                                        IE.pdf.

Cordier:2001:IDB

[207] Frédéric Cordier, Pascal Volino, and Nadia Magnenat-Thalmann. Integrating deformations between bodies and clothes. The Journal of Visualization and Computer Animation, 12(1): 45–53, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/82003013/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=82003013&PLACEBO=
                                        IE.pdf.

Magnenat-Thalmann:2001:Ea

[208] Nadia Magnenat-Thalmann and Bernd Eberhardt. Editorial. The Journal of Visualization and Computer Animation, 12(1):i, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/82003012/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=82003012&PLACEBO=
                                        IE.pdf.

Choi:2001:MVH

[209] Soo-Mi Choi and Myoung-Hee Kim. Motion visualization of human left ventricle with a time-varying deformable model for cardiac diagnosis. The Journal of Visualization and Computer Animation, 12(2):55–66, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/

abstract/83502224/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=83502224&PLACEBO=
IE.pdf.

Choe:2001:PDM

[210] Byoungwon Choe, Hanook Lee, and Hyeong-Seok Ko. Performancedriven muscle-based facial animation. The Journal of Visualization and Computer Animation, 12(2): 67–79, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/83502221/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=83502221&PLACEBO=
IE.pdf.

Lee:2001:RWC

[211] Dongkyoo Lee, Hee-Jung Bae, Chang Tae Kim, Dong-Chun Lee, Dae-Hyun Jung, Nam-Kyung Lee, Kyoo-Ho Lee, Nakhoon Baek, J. Won Lee, Kwan Woo Ryu, and James K. Hahn. Reproducing works of Calder. The Journal of Visualization and Computer Animation, 12(2):81–91, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/83502222/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=83502222&PLACEBO=
IE.pdf.

Besuievsky:2001:ARE

[212] Gonzalo Besuievsky and Xavier Pueyo.

Animating radiosity environments through the Multi-Frame Lighting Method. The Journal of Visualization and Computer Animation, 12(2):93–106,

????  2001.  CODEN JVCAEO.

ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/83502223/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=83502223&PLACEBO=
                                        IE.pdf.

Magnenat-Thalmann:2001:Eb

[213] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 12(2):i, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/83502225/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=83502225&PLACEBO=
                                        IE.pdf.

King:2001:PTM

[214] Scott A. King and Richard E. Parent. A 3D parametric tongue model for animated speech. The Journal of Visualization and Computer Animation, 12 (3):107–115, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/85513714/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=85513714&PLACEBO=
                                        IE.pdf.

Kalisiak:2001:GBM

[215] Maciej Kalisiak and Michiel van de Panne. A grasp-based motion planning algorithm for character animation. The Journal of Visualization and Computer Animation, 12(3):117–

129, ????  2001.  CODEN JV-

CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/85513712/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=85513712&PLACEBO=
IE.pdf.

Skapin:2001:UCP

[216] X. Skapin and P. Lienhardt. Using Cartesian product for animation. The Journal of Visualization and Computer Animation, 12(3):131– 144, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/85513709/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=85513709&PLACEBO=
IE.pdf.

Ehmann:2001:TES

[217] Stephen A. Ehmann, Arthur D. Gregory, and Ming C. Lin. A touch-enabled system for multi-resolution modeling and 3D painting. The Journal of Visualization and Computer Animation, 12 (3):145–157, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/85513715/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=85513715&PLACEBO=
IE.pdf.

Oh:2001:MCE

[218] Kyoung-Su Oh, Byeong-Seok Shin, and Yeong Gil Shin. Mobility culling: an efficient rendering algorithm using temporal coherence. The Journal of Visualization and Computer Animation, 12 (3):159–166, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/85513713/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=85513713&PLACEBO=
                                        IE.pdf.

Handa:2001:VTV

[219] Susumu Handa, Hiroshi Kashiwagi, and Toshikazu Takada. Visualization techniques for 3D vector fields: an application to electrostatic fields of molecules. The Journal of Visualization and Computer Animation, 12(3): 167–180, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/85513710/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=85513710&PLACEBO=
                                        IE.pdf.

Magnenat-Thalmann:2001:Ec

[220] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 12(3):i, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/85513711/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=85513711&PLACEBO=
                                        IE.pdf.

Akleman:2001:FBF

[221] Ergun Akleman, Zeki Melek, and Jeff S. Haberl. Function-based flow modeling and animation. The Journal of Visualization and Computer Animation, 12 (4):181–189, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/89011185/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=89011185&PLACEBO=
IE.pdf.

Benes:2001:UPT

[222] Bedřich Beneš and Enrique Espinosa.

Using particles for 3D texture sculpting. The Journal of Visualization and Computer Animation, 12(4):191– 201, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/89011189/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=89011189&PLACEBO=
IE.pdf.

Hughes:2001:CTT

[223] Charles E. Hughes, J. Michael Moshell, Dean Reed, Diane Z. Chase, and Arlen F. Chase. The Caracol Time Travel Project. The Journal of Visualization and Computer Animation, 12 (4):203–214, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/89011188/START;   http:/
/www3.interscience.wiley.com/cgi-
bin/fulltext?ID=89011188&PLACEBO=
IE.pdf.

Rudomin:2001:MGU

[224] Isaac Rudomín, Roberto Pérez-Urbiola, Maria Elena Melón, and Jose Luis Castillo. Multilayer garments using isosurfaces and physics. The Journal of Visualization and Computer Animation, 12(4):215–226, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/89011187/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=89011187&PLACEBO=
                                        IE.pdf.

Liu:2001:RMA

[225] Zicheng Liu, Zhengyou Zhang, Chuck Jacobs, and Michael Cohen. Rapid modeling of animated faces from video. The Journal of Visualization and Computer Animation, 12(4): 227–240, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/89011186/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=89011186&PLACEBO=
                                        IE.pdf.

Rudomin:2001:E

[226] Isaac Rudomin. Editorial. The Journal of Visualization and Computer Animation, 12(4):i, ???? 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract/89011184/START;   http:/
                                        /www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=89011184&PLACEBO=
                                        IE.pdf.

Grabner:2001:ATI

[227] Markus Grabner. Advanced techniques for interactive visualization of multiresolution meshes. The Journal of Visualization and Computer Animation, 12(5):241–252, December 2001. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bartroli:2001:PSA

[228] A. Vilanova Bartrolí, R. Wegenkittl, and E. Gröller. Projected slabs: approximation of perspective projection and error analysis. The Journal of Visualization and Computer Animation, 12(5): 253–262, December 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bornik:2001:HQT

[229] Alexander Bornik, Konrad Karner, Joachim Bauer, Franz Leberl, and Heinz Mayer. High-quality texture reconstruction from multiple views. The Journal of Visualization and Computer Animation, 12(5):263–276, December 2001. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bittner:2001:ECH

[230] Jiří Bittner and Vlastimil Havran. Exploiting coherence in hierarchical visibility algorithms. The Journal of Visualization and Computer Animation, 12(5): 277–286, December 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

urikovi:2001:GAH

[231] Roman urikovi, Silvester Czanner, and Hirofumi Inoue. Growth animation of human organs. The Journal of Visualization and Computer Animation, 12(5): 287–295, December 2001. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Schmitt:2001:CTB

[232] B. Schmitt, A. Pasko, V. Adzhiev, and C. Schlick. Constructive texturing based on hypervolume modeling. The Journal of Visualization and Computer Animation, 12(5):297–310, December 2001. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

urikovi:2001:EPI

[233] Roman urikovi and Andrej Ferko. Editorial. The Journal of Visualization and Computer Animation, 12(5):iii–iv, December 2001. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Janssen:2002:SISa

[234] J. P. Janssen, M. Egmont-Petersen, E. A. Hendriks, M. J. T. Reinders, R. J. van der Geest, P. C. W. Hogendoorn, and J. H. C. Reiber. Scale-invariant segmentation of dynamic contrast-enhanced perfusion MR images with inherent scale selection. The Journal of Visualization and Computer Animation, 13(1):1–19, February 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Janssen:2002:SISb

[235] J. P. Janssen, M. Egmont-Petersen, E. A. Hendriks, M. J. T. Reinders, R. J. van der Geest, P. C. W. Hogendoorn, and J. H. C. Reiber. Scale-invariant segmentation of dynamic contrast-enhanced perfusion MR images with inherent scale selection. The Journal of Visualization and Computer Animation, 13(1):1–19, February 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bruyns:2002:SIMa

[236] Cynthia D. Bruyns, Steven Senger, Anil Menon, Kevin Montgomery, Simon Wildermuth, and Richard Boyle. A survey of interactive mesh-cutting techniques and a new method for implementing generalized interactive mesh cutting using virtual tools. The Journal of Visualization and Computer Animation, 13(1): 21–42, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bruyns:2002:SIMb

[237] Cynthia D. Bruyns, Steven Senger, Anil Menon, Kevin Montgomery, Simon Wildermuth, and Richard Boyle. A survey of interactive mesh-cutting techniques and a new method for implementing generalized interactive mesh cutting using virtual tools. The Journal of Visualization and Computer Animation, 13(1): 21–42, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Evesque:2002:BASa

[238] F. Evesque, S. Gerlach, and R. D. Hersch. Building 3D anatomical scenes on the Web. The Journal of Visualization and Computer Animation, 13(1): 43–52, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Evesque:2002:BASb

[239] F. Evesque, S. Gerlach, and R. D. Hersch. Building 3D anatomical scenes on the Web. The Journal of Visualization and Computer Animation, 13(1): 43–52, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Snedeker:2002:CHFa

[240] J. G. Snedeker, M. Bajka, J. M. Hug, G. Székely, and P. Niederer. The creation of a high-fidelity finite element model of the kidney for use in trauma research. The Journal of Visualization and Computer Animation, 13(1): 53–64, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Snedeker:2002:CHFb

[241] J. G. Snedeker, M. Bajka, J. M. Hug, G. Székely, and P. Niederer. The creation of a high-fidelity finite element model of the kidney for use in trauma research. The Journal of Visualization and Computer Animation, 13(1): 53–64, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Archip:2002:LMDa

[242] Neculai Archip, Pierre-Jean Erard, Jean-Marie Haefliger, and Jean-Francois Germond. Lung metastasis detection and visualization on CT images: a knowledge-based method. The Journal of Visualization and Computer Animation, 13(1):65–76, February 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Archip:2002:LMDb

[243] Neculai Archip, Pierre-Jean Erard, Jean-Marie Haefliger, and Jean-Francois Germond. Lung metastasis detection and visualization on CT images: a knowledge-based method. The Journal of Visualization and Computer Animation, 13(1):65–76, February 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Langlotz:2002:SAOa

[244] Frank Langlotz. State-of-the-art in orthopaedic surgical navigation with a focus on medical image modalities. The Journal of Visualization and Computer Animation, 13(1):77–83, February 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Langlotz:2002:SAOb

[245] Frank Langlotz. State-of-the-art in orthopaedic surgical navigation with a focus on medical image modalities. The Journal of Visualization and Computer Animation, 13(1):77–83, February 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Harders:2002:NPIa

[246] Matthias Harders, Simon Wildermuth, and Gábor Székely. New paradigms for interactive 3D volume segmentation. The Journal of Visualization and Computer Animation, 13(1):85–95, February 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Harders:2002:NPIb

[247] Matthias Harders, Simon Wildermuth, and Gábor Székely. New paradigms for interactive 3D volume segmentation. The Journal of Visualization and Computer Animation, 13(1):85–95, February 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Magnenat-Thalmann:2002:EPa

[248] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 13(1):i, February 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Picinbono:2002:IRS

[249] Guillaume Picinbono, Jean-Christophe Lombardo, Hervé Delingette, and Nicholas Ayache. Improving realism of a surgery simulator: linear anisotropic elasticity, complex interactions and force extrapolation. The Journal of Visualization and Computer Animation, 13(3): 147–167, July 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pal:2002:LMA

[250] Prabir K. Pal and Biswajit Sarkar. A language to model animation out of behaviour-embedded graphical components. The Journal of Visualization and Computer Animation, 13(3):169– 185, July 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yang:2002:NPM

[251] Shu-Kai Yang, Chin-Chen Chang, Ding-Zhou Duan, and Ming-Fen Lin. A novel progressive modelling algorithm for 3D models. The Journal of Visualization and Computer Animation, 13(3):187– 198, July 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Magnenat-Thalmann:2002:EPc

[252] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 13(3):i, July 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pollefeys:2002:VMI

[253] Marc Pollefeys and Luc Van Gool. Visual modelling: from images to images. The Journal of Visualization and Computer Animation, 13(4):199– 209, September 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Han:2002:PFM

[254] Mei Han and Takeo Kanade. A perspective factorization method for Euclidean reconstruction with uncalibrated cameras. The Journal of Visualization and Computer Animation, 13(4): 211–223, September 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zhang:2002:SVM

[255] Li Zhang, Guillaume Dugas-Phocion, Jean-Sebastien Samson, and Steven M. Seitz. Single-view modelling of freeform scenes. The Journal of Visualization and Computer Animation, 13(4): 225–235, September 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Hung:2002:APO

[256] Yi-Ping Hung, Chu-Song Chen, Yu-Pao Tsai, and Szu-Wei Lin. Augmenting panoramas with object movies by generating novel views with disparity-based view morphing. The Journal of Visualization and Computer Animation, 13(4): 237–247, September 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Tong:2002:LLL

[257] Xin Tong, Jinxiang Chai, and Heung-Yeung Shum. Layered lumigraph with LOD control. The Journal of Visualization and Computer Animation, 13(4):

249–261, September 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Ng:2002:EPI

[258] Teck Khim Ng and Tat-Seng Chua. Editorial. The Journal of Visualization and Computer Animation, 13(4):iii–iv, September 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Devillers:2002:PEB

[259] Frédéric Devillers, Stéphane Donikian, Fabrice Lamarche, and Jean-François Taille. A programming environment for behavioural animation. The Journal of Visualization and Computer Animation, 13(5):263–274, December 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zhongxiang:2002:MAC

[260] Luo Zhongxiang, Zhuang Yueting, Liu Feng, and Pan Yunhe. Multiple animated characters motion fusion. The Journal of Visualization and Computer Animation, 13(5):275–285, December 2002. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Gillies:2002:EMA

[261] M. F. P. Gillies and N. A. Dodgson. Eye movements and attention for behavioural animation. The Journal of Visualization and Computer Animation, 13(5):287–300, December 2002. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pelachaud:2002:SFE

[262] Catherine Pelachaud and Isabella Poggi. Subtleties of facial expressions in embodied agents. The Journal of Visualization and Computer Animation, 13(5): 301–312, December 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bailenson:2002:GTP

[263] Jeremy N. Bailenson, Andrew C. Beall, and Jim Blascovich. Gaze and task performance in shared virtual environments. The Journal of Visualization and Computer Animation, 13(5):313– 320, December 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Ballin:2002:EPI

[264] Daniel Ballin, Jeff Rickel, and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 13(5):i–ii, December 2002. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Min:2003:ASD

[265] KyungHa Min, In-Kwon Lee, and Chan-Mo Park. Adaptive space decomposition for fast visualization of soft objects. The Journal of Visualization and Computer Animation, 14(1):1–19, February 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yanyun:2003:MRS

[266] Chen Yanyun, Hanqiu Sun, Lin Hui, and Enhua Wu. Modelling and rendering of snowy natural scenery using multi-mapping techniques. The Journal of Visualization and Computer Animation, 14(1):21–30, February 2003. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Faraway:2003:RMM

[267] Julian Faraway. Regression modeling of motion with endpoint constraints. The Journal of Visualization and Computer Animation, 14(1):31–41, February 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Baranoski:2003:SA

[268] Gladimir V. G. Baranoski, Jon G. Rokne, Peter Shirley, Trond S. Trondsen, and Rui Bastos. Simulating the aurora. The Journal of Visualization and Computer Animation, 14(1):43–59, February 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Magnenat-Thalmann:2003:EPI

[269] Nadia Magnenat-Thalmann and Daniël Thalmann. Editorial. The Journal of Visualization and Computer Animation, 14(1):iii, February 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Chen:2003:RCF

[270] Haixin Chen, Jürgen Hesser, and Reinhard Männer. Ray casting free-form deformed-volume objects. The Journal of Visualization and Computer Animation, 14(2):61–72, May 2003. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Hase:2003:HGS

[271] Kazunori Hase, Kazuo Miyashita, Sooyol Ok, and Yoshiki Arakawa. Human gait simulation with a neuromusculoskeletal model and evolutionary computation. The Journal of Visualization and Computer Animation, 14(2):

73–92, May 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Park:2003:SAD

[272] In Kyu Park, Sang Wook Lee, and Sang Uk Lee. Shape-adaptive 3-D mesh simplification based on local optimality measurement. The Journal of Visualization and Computer Animation, 14(2): 93–109, May 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Magnenat-Thalmann:2003:EPa

[273] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 14(2):i, May 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kampel:2003:APA

[274] Martin Kampel and Robert Sablatnig.

An automated pottery archival and reconstruction system. The Journal of Visualization and Computer Animation, 14(3):111–120, July 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Taylor:2003:NIT

[275] J. Taylor, J.-A. Beraldin, G. Godin, L. Cournoyer, R. Baribeau, F. Blais, M. Rioux, and J. Domey. NRC 3D imaging technology for museum and heritage applications. The Journal of Visualization and Computer Animation, 14(3): 121–138, July 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yang:2003:RDS

[276] Hyun S. Yang, Taewoo Han, and Juho Lee. Reconstruction of dynamic 3D scene based on visual hull and view morphing. The Journal of Visualization and Computer Animation, 14(3): 139–148, July 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Cappellini:2003:AAO

[277] V. Cappellini, M. Barni, M. Corsini, A. De Rosa, and A. Piva. ArtShop: an art-oriented image-processing tool for cultural heritage applications. The Journal of Visualization and Computer Animation, 14(3):149–158, July 2003. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Bakker:2003:TCD

[278] Geeske Bakker, Frans Meulenberg, and Jan de Rode. Truth and credibility as a double ambition: reconstruction of the built past, experiences and dilemmas. The Journal of Visualization and Computer Animation, 14(3):159–167, July 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Magnenat-Thalmann:2003:EPb

[279] Nadia Magnenat-Thalmann and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 14(3):i, July 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Fan:2003:IFC

[280] Zhaowei Fan, Huagen Wan, and Shuming Gao. IBCD: a fast collision detection algorithm based on image space using OBB. The Journal of Visualization and Computer Animation, 14(4): 169–181, September 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Masuda:2003:SDV

[281] Tomohito Masuda, Setsuo Imazu, Supatana Auethavekiat, Tsuyoshi Furuya, Kunihiko Kawakami, and Katsushi Ikeuchi. Shape difference visualization for ancient bronze mirrors through 3D range images. The Journal of Visualization and Computer Animation, 14(4): 183–196, September 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Moriya:2003:IMI

[282] Toshio Moriya, Kei Utsugi, Fumiko Beniyama, and Haruo Takeda. Image modification for immersive projection display based on pseudo-projection models. The Journal of Visualization and Computer Animation, 14(4):197– 210, September 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Zhang:2003:TCR

[283] Mingmin Zhang, Zhigeng Pan, and Pheng-Ann Heng. Time-critical rendering algorithm with incorporation of LoD, visibility culling and object impostor. The Journal of Visualization and Computer Animation, 14(4):211– 223, September 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Pletinckx:2003:HPT

[284] Daniel Pletinckx, Neil Silberman, and Dirk Callebaut. Heritage presentation through interactive storytelling: a new multimedia database approach. The Journal of Visualization and Computer Animation, 14(4):225–231, September 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Burdea:2003:EP

[285] Greg Burdea and Daniel Thalmann. Editorial. The Journal of Visualization and Computer Animation, 14(5):i, December 2003. CODEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Rutten:2003:IPS

[286] Anja Rutten, Sue Cobb, Helen Neale, Steven Kerr, Anne Leonard, Sarah Parsons, and Peter Mitchell. The AS interactive project: single-user and collaborative virtual environments for people with high-functioning autistic spectrum disorders. The Journal of Visualization and Computer Animation, 14(5): 233–241, December 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Yano:2003:DGR

[287] Hiroaki Yano, Kaori Kasai, Hideyuki Saitou, and Hiroo Iwata. Development of a gait rehabilitation system using a locomotion interface. The Journal of Visualization and Computer Animation, 14(5):243–252, December 2003. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Martin:2003:EOC

[288] Olivier Martin, Benjamin Julian, Laurence Boissieux, Jean-Dominique Gascuel, and Claude Prablanc. Evaluating online control of goal-directed arm movement while standing in virtual visual environment. The Journal of Visualization and Computer Animation, 14(5):253–260, December 2003. CO-DEN JVCAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Kizony:2003:AIV

[289] Rachel Kizony, Noomi Katz, and Patrice L. (Tamar) Weiss. Adapting an immersive virtual reality system for rehabilitation. The Journal of Visualization and Computer Animation, 14(5): 261–268, December 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).

Camurri:2003:AMT

[290] Antonio Camurri, Barbara Mazzarino, Gualtiero Volpe, Pietro Morasso, Federica Priano, and Cristina Re. Application of multimedia techniques in the physical rehabilitation of Parkinson’s patients. The Journal of Visualization and Computer Animation, 14(5): 269–278, December 2003. CODEN JV-CAEO. ISSN 1049-8907 (print), 1099-1778 (electronic).